a motel charges 80.00 per day for double occupancy and 70.00 per day for single occupancy. If 80 rooms are occupied one night for a total of 5800.00 how many rooms of each kind are occupied?
Answer by rfer(16322)

d+s=80, d=80-s
80d+70s=5800
80(80-s)+70s=5800
6400-80s+70s=5800
-10s=-600
s=60
d=20